OpenStudy (anonymous):

What is the simplified form of x plus 1 over x squared plus x minus 6 divided by x squared plus 5x plus 4 over x plus 3 ?

OpenStudy (zehanz):

If I write it in the equation editor, it is simpler already ;)\[\frac{ x+1 }{ x^2 }+\frac{x-6}{x^2}+\frac{5x+4}{x+3}\] You can start with the two leftmost fractions; they have the same denominator, so can be written as one fraction.

OpenStudy (anonymous):

Which makes it \[\frac{ 2x-5 }{ x ^{2} }+\frac{ 5x+4 }{ x+3 }\] right?

OpenStudy (zehanz):

Yes! Now you have write the fractions with a new denominator, that is the same for both. Do you know how to do that?

OpenStudy (zehanz):

Hint: use x²(x+3) as the new denominator.

OpenStudy (anonymous):

\[\frac{ 2x-5 }{ x ^{3} +3x^{2} }+\frac{ 5x+4 }{ x ^{3}+3x ^{2} }\] right?

OpenStudy (zehanz):

Right, although I would prefer to leave the denominator as x²(x+3). Now write it as one fraction and simplify the numerator.

OpenStudy (anonymous):

\[\frac{ 7x-1 }{ x ^{2}(x+3) }\] right?

OpenStudy (zehanz):

Oops, I only looked at the denominator. YOu can't just change them without changing the numerator as well:\[\frac{ (2x-5)(x+3) }{ x^2(x+3) }+\frac{ x^2(5x+4) }{ x^2(x+3) }\]

OpenStudy (zehanz):

Denominator of first fraction has been multiplied with x+3, so do that in the numerator as well. Second one: both multiplied with x²

OpenStudy (zehanz):

Have to go now, back in half an hour...

OpenStudy (anonymous):

Okay, so then it's \[\frac{ 2x ^{3}+x-15 }{ x ^{2}(x+3) }+\frac{ 5x^{3}+4x^{2} }{ x^{2}(x+3) }\]Right?

OpenStudy (anonymous):

Okay.

OpenStudy (zehanz):

Hey that's right! Now write it as one fraction. (now I really must go, sorry..)

OpenStudy (anonymous):

Okay, see you later!

OpenStudy (anonymous):

\[\frac{ 7x^{3}+4x^{2}+x-15 }{ x^{2}(x+3) }\]right?

OpenStudy (zehanz):

@HarvestMan: You have got it right!

OpenStudy (anonymous):

Awesome!! Can it be simplified any further?

OpenStudy (anonymous):

The only answers for my homework are 1 over the quantity x plus 3 times the quantity x plus 4 1 over the quantity x plus 3 times the quantity x minus 2 1 over the quantity x plus 4 times the quantity x minus 2 or 1 over the quantity x plus 3 times the quantity x plus 1

OpenStudy (zehanz):

No, because this could only be done if you'd factor the numerator. If there were common factors, the would cancel. But: x is not a common factor, and x-3 neither (to see why, set x=3 in the numerator - it won't come out at 0), so there is nothing to be done about it...

OpenStudy (zehanz):

If these are your only options, then have we worked with the right fractions from the beginning?

OpenStudy (anonymous):

Yes, we have. But, can we simplify our answer any further?

OpenStudy (anonymous):

oh

OpenStudy (zehanz):

Hold on, I'll check everything from the beginning....

OpenStudy (anonymous):

Okay

OpenStudy (zehanz):

When I do the last step again, I get:

OpenStudy (zehanz):

\[\frac{ (2x-5)(x+3) +x^2(5x+4)}{ x^2(x+3) }=\frac{ 2x^2+6x-5x-15+5x^3+4x^2 }{ x^2(x+3) }=\]\[\frac{ 5x^3+6x^2+x-15 }{ x^2(x+3) }\]Cannot simplify this further...

OpenStudy (anonymous):

Hm.. which one of the answers do you think would be the closest to this?

OpenStudy (anonymous):

I'm guessing answer 1 is the closest

OpenStudy (anonymous):

I've got to go, talk to you later!

OpenStudy (zehanz):

OK, wouldn't have a clue, I'm afraid...

OpenStudy (anonymous):

I'll just go with answer #1